# Closing the Data-Efficiency Gap Between Autoregressive and Masked Diffusion LLMs

### v1

@@ Abstract
Despite autoregressive large language models (arLLMs) being the current dominant paradigm in language modeling, they resist knowledge injection via fine-tuning due to inherent shortcomings such as the “reversal curse” — the challenge of answering questions that reverse the original information order in the training sample. Masked diffusion large language models (dLLMs) are rapidly emerging as a powerful alternative to the arLLM paradigm, with evidence of better data efficiency and free of the “reversal curse” in pre-training. However, it is unknown whether these advantages extend to the post-training phase, i.e. whether pre-trained dLLMs can easily acquire new knowledge through fine-tuning. On three diverse datasets, we fine-tune arLLMs and dLLMs, evaluating them with forward and backward style Question Answering (QA) to probe knowledge generalization and the reversal curse. Our results confirm that arLLMs critically rely on extensive data augmentation via paraphrases for QA generalization, and paraphrases are only effective when their information order matches the QA style. Conversely, dLLMs achieve high accuracies on both forward and backward QAs without paraphrases; adding paraphrases yields only marginal gains. Lastly, inspired by the dLLM’s performance, we introduce a novel masked fine-tuning paradigm for knowledge injection into pre-trained arLLMs. This proposed method successfully and drastically improves the data efficiency of arLLM fine-tuning, effectively closing the performance gap with dLLMs.
@@ 1 Introduction
Despite auto-regressive large language models (arLLMs) being the main contributors to the modern success of language modeling, studies have demonstrated the difficulty of injecting new knowledge into pre-trained arLLMs by fine-tuning on documents that are not in the pre-training dataset (Ovadia et al., 2023; Mecklenburg et al., 2024; Gekhman et al., 2024; Soudani et al., 2024; Zhao et al., 2025; Lampinen et al., 2025). Fine-tuned models typically generalize poorly to downstream tasks such as question-answering (QA). An example failure mode is the famous “reversal curse”, that LLMs fail to answer the questions in the reversed order of the training text (Berglund et al., 2023). Fine-tuning on multiple rewrites (i.e., paraphrases) of the documents can mitigate generalization issues but still falls behind in-context learning based external memory systems like RAG (Ovadia et al., 2023; Mecklenburg et al., 2024). This pitfall of arLLMs is a major obstacle that limits current models from being flexible life-long learners via weight updates.
As alternatives to the auto-regressive models, several recent masked diffusion large language models (dLLMs) have been scaled up to be as capable as arLLMs on multiple downstream tasks, with additional advantages such as high-throughput decoding of multiple tokens simultaneously (Nie et al., 2025a; b; Ye et al., 2025). Instead of factorizing the joint sequence probability strictly left-to-right, dLLMs learn the factorization under many permutations of token positions, enabling them to condition on any subset of tokens and predict the remainder. Though such an objective is harder than learning autoregressive factorization (Kim et al., 2025), dLLMs do not suffer from “reversal curse(Nie et al., 2025b), and can achieve lower validation loss (i.e. token cross-entropy loss) than arLLMs in data-constrained settings (Prabhudesai et al., 2025; Ni & the team, 2025). However, most of the dLLMs studies have focused on the properties in the pre-training phase. Little is known about whether dLLMs also have advantages in the post-training phase, such as knowledge injection by fine-tuning.
In this study, we use three datasets to compare the data efficiency and performance of knowledge injection by fine-tuning in arLLM and dLLM models. We also introduce a novel masked fine-tuning paradigm for arLLMs that emulates a diffusion-style mask reconstruction loss. Across all datasets, dLLMs show a consistent data-efficiency advantage over arLLMs, and our masked fine-tuning largely closes this gap, bringing arLLMs to strong performance without relying on paraphrases. More specifically, we show the following results:
@@
  • We propose a masked fine-tuning paradigm that fine-tunes arLLMs in a “masked diffusion” way by giving masked samples in the context with instructions to recover the mask, and set the unmasked sample as the supervised fine-tuning target. The novel method closes the performance gap between arLLMs and dLLMs fine-tuning: it achieves strong performance in both forward and backward questions without paraphrases.
@@ 2.1 Knowledge injection by fine-tuning and Reversal curse
An AI system intended for real-world deployment should continually acquire and integrate new information to adapt to evolving environments. Though LLMs have been successful on numerous tasks, they struggle to incorporate new knowledge into their weights. At least two factors contribute to this challenge. First, fine-tuning on new tasks can induce catastrophic forgetting of previously learned capabilities (Luo et al., 2023; Wang et al., 2023; Zhai et al., 2023; Zhang & Wu, 2024; Chen et al., 2024; Ren et al., 2024). Second, fine-tuning primarily reshapes surface behavior (e.g., tone and format) without reliably integrating new factual knowledge (Ovadia et al., 2023; Mecklenburg et al., 2024; Gekhman et al., 2024; Soudani et al., 2024; Zhao et al., 2025; Lampinen et al., 2025).
A failure mode of learning knowledge in the text is the “reversal curse”, that after learning statements of the form “A is B”, the model does not generalize to its inverse form “B is A”. The reversal curse has been observed across the training phases and models (Berglund et al., 2023; Allen-Zhu & Li, 2025; Lv et al., 2024; Lin et al., 2024; Guo et al., 2024; Golovneva et al., 2024; Lu et al., 2024). Even strong commercial models like GPT-4 and GPT-4o show signs of the reversal curse (Berglund et al., 2023; Nie et al., 2025b). The cause of the reversal curse has been theoretically attributed to an inherent limitation of the autoregressive training objective (Zhu et al., 2024; Kitouni et al., 2024) (See Appendix A.6). Common approaches to mitigate the reversal curse in autoregressive models include: (i) augmenting the training set with paraphrases (Lu et al., 2024), which requires substantial computation to construct; (ii) augmenting the training set with reordered sequences (Guo et al., 2024; Golovneva et al., 2024), which often violate natural-language grammar and degrade overall language modeling performance; and (iii) replacing causal attention with bidirectional attention (Lv et al., 2024), which struggles to retrieve long information (e.g., a person’s description). In contrast, our masked fine-tuning paradigm for arLLMs addresses the reversal curse without constructing paraphrase augmentations or altering the autoregressive objective. With only adjustments to the fine-tuning query and keeping the original documents as optimization targets, our method attains high accuracy across all datasets we evaluated.
@@ 2.2 Masked diffusion language models
Recently, dLLMs have emerged as a strong competitor to arLLMs (Sahoo et al., 2024; Nie et al., 2025b; Ye et al., 2025). Comparing to autoregressive models, dLLMs use encoder-only transformers to generate text by iteratively unmasking tokens via a reversed discrete diffusion process. The training objective is to minimize the mask reconstruction loss Nie et al. (2025b):
@@
$\bm{x_{0}}$ is an original sequence of length $L$ sampled from the training data. The masking process is governed by the mask ratio $t$, which is sampled uniformly, resulting in the corrupted sequence $\bm{x_{t}}$. The set $\mathbf{M}$ denotes the indices of the tokens that were masked by the forward process at ratio $t$. The $\ell$-th token is considered for the loss only if it was masked. Such a loss objective has been shown to be the negative evidence lower bound (ELBO) on the data likelihood (Shi et al., 2024).
Recent studies report dLLMs are more data-efficient than arLLMs. When the training data is scarce, dLLMs keep improving with repeated use of the data and surpass arLLMs on validation loss, while arLLMs saturate the validation loss or increase it due to overfitting (Prabhudesai et al., 2025; Ni & the team, 2025). Prabhudesai et al. (2025) further shows that the lower validation loss in dLLMs can generalize to downstream tasks like ARC-Easy, and attributes its data efficiency to random masks as implicit data augmentation. However, whether these advantages persist in new knowledge acquisition during post-training—where the model needs to learn generalizable knowledge through small fine-tuning sets—remains unclear.
2.3 Change order training of arLLM
In the following sections, we introduce a novel masked fine-tuning paradigm for arLLMs that is distinct from prior studies employing permuted-order training. Specifically, XLNet (Yang et al., 2019) proposed bidirectional context learning using Permutation Language Modeling, yet this required engineering a specialized Two-Stream Self-Attention architecture. Subsequent work like MAC (Shih et al., 2022) focused on optimizing the training efficiency of these any-order models but similarly involved modifications to the underlying attention mechanism. Furthermore, methods like the fill-in-the-middle objective (Bavarian et al., 2022) functioned solely as a pre-training objective for infilling, without addressing the post-training knowledge acquisition challenges in downstream tasks such as QA. In contrast, our proposed masked fine-tuning paradigm focuses exclusively on the post-training phase, effectively capturing the benefits of bidirectional context to improve knowledge generalization without necessitating any alteration to the core pre-trained autoregressive architecture.
@@ 3 Datasets and experimental setups
We focus on assessing LLMs’ ability to learn new knowledge through fine-tuning. More specifically, LLMs are fine-tuned on a set of documents that contain knowledge unknown to the base LLM, and evaluated by open-ended QA tasks. The correctness of an answer is evaluated by the well-adopted ROUGE-1 score (Lin et al., 2024; Jiang et al., 2025) between the generated answer and the ground truth answer, which we report as “accuracy.” It measures the proportion of the words in the ground truth answer that appear in the generated answer. To better demonstrate the generation quality, we also show examples of model responses in all the experiments in A.5.
@@
We use three representative datasets. Two are existing synthetic datasets from previous studies on the reversal curse; one is constructed using recent Wikipedia articles. Each dataset has been augmented with paraphrases. See examples of each dataset in Appendix A.3.
@@
The NameDescription (Berglund et al., 2023) contains 60 statements of different fictitious individuals, 30 each of the form “[name] is [description]” (N2D) and “[description] is [name]” (D2N). Lin et al. (2024) extended the dataset with an open-ended QA testing set. For each type of statement, the QA set contains two types of questions: “What is the name related to a given description” and “What is the description of a given name”. Depending on whether the question is aligned with the original statement, each question is classified as “forward” or “backward” question (e.g. N2D statement with “What is the description of a given name” type of question is a forward question). The dataset also contains a paraphrase set, in which each statement is rewritten into 30 different versions, but the order of [name] and [description] in the paraphrases is always preserved as in the original statement (either N2D or D2N).
@@
The Biography dataset is proposed in Allen-Zhu & Li (2024; 2025). Since the original dataset is not publicly available, we used a subset of 100 samples from a replication (Zheng et al., 2025). Each sample is a 6-sentence paragraph about a fictitious individual, detailing their birth city, birthday, college, and job information. Note that the name only appears in the first sentence and is replaced with a pronoun in the following sentences; thus, questions about the name are considered backward questions. Each sample also includes a paraphrase set of 5 paraphrases; the paraphrases do not change the order of the sentences but only alter the wording while preserving the information. The testing QA set has both forward (i.e., asking for an attribute given the name) and backward styles (i.e., asking for the name given 3 attributes from the person) questions.
@@
The Wiki dataset contains 92 Wikipedia articles, constructed according to the protocol described in Pan et al. (2025). We crawl the Wikipedia pages under the category “2025 by month”, and then further filter out the pages that were created before the year 2025. This procedure ensures that these real-world events are recent enough that both dLLM and arLLM models are not aware of them, which is justified by the models’ accuracy before fine-tuning (Table 3). For each wiki article, we use GPT-o3-mini to generate QA pairs in both forward and backward styles. By prompting GPT-o3-mini, we construct two different paraphrase sets: one that retains the information in place while only changing the wording (same-order paraphrases); the other also changes the order of information in the article (permute-order paraphrases). 10 paraphrases of each type are generated for every wiki article. More details on constructing the datasets are provided in Appendix A.3.
@@
We chose Llama-3.1-8B-Instruct (Dubey et al., 2024) and LLaDA-8B-Instruct (Nie et al., 2025b) models as representatives of arLLM and dLLM to conduct the experiments, as they perform similarly on the benchmarks and are of comparable sizes. Fine-tuning and evaluation configurations are provided in the Appendix A.4.
NameDescription Biography
N2D-fwd N2D-bwd D2N-fwd D2N-bwd Fwd Bwd
arLLM before fine-tuning 0.072 0.000 0.054 0.000 0.001 0.000
arLLM w/o paraphrases 0.374 0.000 0.017 0.027 0.121 0.002
arLLM w paraphrases 0.910 0.004 0.925 0.071 0.962 0.001
Table 1: Fine-tuning performance of arLLM on the NameDescription and Biography datasets.
@@
arLLM before fine-tuning 0.164 0.127
arLLM w/o paraphrases 0.377 0.282
arLLM w same-order paraphrases 0.685 0.396
arLLM w permute-order paraphrases 0.721 0.628
@@
Table 2: Fine-tuning performance of arLLM on the Wiki dataset.
4 arLLM knowledge injection relies on paraphrases
We first show that knowledge injection by fine-tuning in arLLMs heavily relies on paraphrases. This is known in previous studies (Berglund et al., 2023; Allen-Zhu & Li, 2025; Lin et al., 2024; Guo et al., 2024; Golovneva et al., 2024). We demonstrate this observation on three datasets to set baselines for comparison with dLLM and our novel paradigm in the following sections.
We fine-tune Llama-3.1-8B-Instruct on dataset samples using the pre-training format (i.e., without a chat template). Without paraphrases, backward accuracy on the NameDescription and Biography datasets is close to 0, while the forward accuracy of NameDescription N2D and Biography does not completely fail but is still poor (Table 1). Adding paraphrases drastically raises forward accuracy close to 1, while backward accuracy remains close to 0. Paraphrases do not help backward accuracy in NameDescription and Biography datasets because the construction of these datasets does not change the semantic order of the sentences. The trend is similar in the Wiki dataset (Table 2). While the same-order paraphrases significantly increase forward accuracy, they only mildly increase backward accuracy. Using permute-order paraphrases increases both forward and backward accuracy, and the gap between them is smaller. Note that, due to the naturalness of this dataset, pre–fine-tuning accuracies are not as close to zero as in the other datasets (Table 2); nonetheless, they are sufficiently low to demonstrate the effectiveness of fine-tuning.
These results suggest that, in arLLM fine-tuning, paraphrases significantly improve QA accuracy, but help backward questions only when the paraphrases change the information order in the sentences to be more aligned with the backward style. Note that the accuracy difference between fine-tuning with paraphrases and without paraphrases is not due to different training steps; in both cases, we train the models with sufficiently large epoch numbers; the reported accuracy is from the best checkpoints during the training (Figure 2, Appendix 6).
5 dLLM knowledge injection
NameDescription Biography Wiki
N2D-fwd N2D-bwd D2N-fwd D2N-bwd Fwd Bwd Fwd Bwd
arLLM before fine-tuning 0.072 0.000 0.054 0.000 0.001 0.000 0.164 0.127
dLLM before fine-tuning 0.030 0.000 0.028 0.000 0.030 0.000 0.210 0.156
arLLM w/o paraphrases 0.374 0.000 0.017 0.027 0.121 0.002 0.377 0.282
arLLM w paraphrases 0.910 0.004 0.925 0.071 0.962 0.001 0.685 0.396
dLLM w/o paraphrases 0.873 0.913 0.864 0.790 0.892 0.696 0.908 0.778
dLLM w paraphrases 0.967 0.994 0.994 0.973 0.991 0.857 0.900 0.785
Masked arLLM w/o paraphrases 0.658 0.949 0.992 0.923 0.971 0.598 0.980 0.930
Masked arLLM w paraphrases 0.969 0.996 0.928 0.832 0.965 0.816 0.905 0.794
Table 3: Fine-tuning performance of arLLM, dLLM and masked arLLM on all three datasets. The paraphrases used for the Wiki dataset are the same-order paraphrase set.

### v2

@@ Abstract
Despite autoregressive large language models (arLLMs) being the current dominant paradigm in language modeling, effectively updating these models to incorporate new factual knowledge still remains difficult. They resist knowledge injection via fine-tuning due to inherent shortcomings such as the “reversal curse” — the challenge of answering questions that reverse the original information order in the training sample. Masked diffusion large language models (dLLMs) are rapidly emerging as a powerful alternative to the arLLM paradigm, with evidence of better data efficiency and free of the “reversal curse” in pre-training. However, it is unknown whether these advantages extend to the post-training phase, i.e. whether pre-trained dLLMs can easily acquire new knowledge through fine-tuning. On three diverse datasets, we fine-tune arLLMs and dLLMs, evaluating them with forward and backward style Question Answering (QA) to probe knowledge generalization and the reversal curse. Our results confirm that arLLMs critically rely on extensive data augmentation via paraphrases for QA generalization, and paraphrases are only effective when their information order matches the QA style. Conversely, dLLMs achieve high accuracies on both forward and backward QAs without paraphrases; adding paraphrases yields only marginal gains. Inspired by the dLLM’s performance, we introduce a novel masked fine-tuning paradigm for knowledge injection into pre-trained arLLMs. This proposed method successfully and drastically improves the data efficiency of arLLM fine-tuning, effectively closing its performance gap with dLLMs. We further show that the masked fine-tuning paradigm of arLLMs can be extended to the supervised fine-tuning (SFT) of mathematical capability. Across two models and two datasets, our masked SFT outperforms regular SFT.
@@ 1 Introduction
Despite auto-regressive large language models (arLLMs) being the main contributors to the modern success of language modeling, studies have demonstrated the difficulty of injecting new knowledge into pre-trained arLLMs by fine-tuning on documents that are not in the pre-training dataset (ovadia2023fine; mecklenburg2024injecting; gekhman2024does; soudani2024fine; zhao2025style; lampinen2025generalization). Fine-tuned models typically generalize poorly to downstream tasks such as question-answering (QA). An example failure mode is the famous “reversal curse”, that LLMs fail to answer the questions in the reversed order of the training text (berglund2023reversal). Fine-tuning on multiple rewrites (i.e., paraphrases) of the documents can mitigate generalization issues but still falls behind in-context learning based external memory systems like RAG (ovadia2023fine; mecklenburg2024injecting). This pitfall of arLLMs is a major obstacle that limits current models from being flexible life-long learners via weight updates.
As alternatives to the auto-regressive models, several recent masked diffusion large language models (dLLMs) have been scaled up to be as capable as arLLMs on multiple downstream tasks, with additional advantages such as high-throughput decoding of multiple tokens simultaneously (ICLR2025_ce1c1ff5; llada; ye2025dream). Instead of factorizing the joint sequence probability strictly left-to-right, dLLMs learn the factorization under many permutations of token positions, enabling them to condition on any subset of tokens and predict the remainder. Though such an objective is harder than learning autoregressive factorization (kim2025train), dLLMs do not suffer from “reversal curse(llada), and can achieve lower validation loss (i.e. token cross-entropy loss) than arLLMs in data-constrained settings (prabhudesai2025diffusion; ni2025difflm). However, most of the dLLMs studies have focused on the properties in the pre-training phase. Little is known about whether dLLMs also have advantages in the post-training phase, such as knowledge injection by fine-tuning.
In this study, we use three datasets to compare the data efficiency and performance of knowledge injection by fine-tuning in arLLM and dLLM models. We refer to “data-efficiency” as how much fine-tuning data (including paraphrases) is required for a pre-trained model to generalize the knowledge in the data to downstream tasks such as QA. We introduce a novel masked fine-tuning paradigm for arLLMs that emulates a diffusion-style mask reconstruction loss. Across all datasets, dLLMs show a consistent data-efficiency advantage over arLLMs, and our masked fine-tuning largely closes this gap, bringing arLLMs to strong performance without relying on paraphrases. We further show that adapting demasking objection to supervised fine-tuning (SFT) of arLLM outperforms regular SFT on math tasks. More specifically, we show the following results:
@@
  • We propose a masked fine-tuning paradigm that fine-tunes arLLMs in a “masked infilling” way by giving masked samples in the context with instructions to recover the mask, and set the unmasked sample as the supervised fine-tuning target. The novel method closes the performance gap between arLLMs and dLLMs fine-tuning: it allows arLLMs of different sizes and model families to achieve strong performance in both forward and backward questions without paraphrases.
  • We extend the masked fine-tuning paradigm to SFT tasks, and show that across two arLLMs and two math datasets masked SFT outperforms regular SFT.
@@ 2.1 Knowledge injection by fine-tuning and Reversal curse
An AI system intended for real-world deployment should continually acquire and integrate new information to adapt to evolving environments. Though LLMs have been successful on numerous tasks, they struggle to incorporate new knowledge into their weights. At least two factors contribute to this challenge. First, fine-tuning on new tasks can induce catastrophic forgetting of previously learned capabilities (luo2023empirical; wang2023trace; zhai2023investigating; zhang2024dissecting; chen2024chatgpt; ren2024analyzing). Second, fine-tuning primarily reshapes surface behavior (e.g., tone and format) without reliably integrating new factual knowledge (ovadia2023fine; mecklenburg2024injecting; gekhman2024does; soudani2024fine; zhao2025style; lampinen2025generalization).
A failure mode of learning knowledge in the text is the “reversal curse”, that after learning statements of the form “A is B”, the model does not generalize to its inverse form “B is A”. The reversal curse has been observed across the training phases and models (berglund2023reversal; zhu24b; lv-etal-2024-analysis; lin2024delving; guo-etal-2024-mitigating; golovneva2024reverse; lu2024rethinking). Even strong commercial models like GPT-4 and GPT-4o show signs of the reversal curse (berglund2023reversal; llada). The cause of the reversal curse has been theoretically attributed to an inherent limitation of the autoregressive training objective (zhu2024towards; kitouni2024factorization) (more discussion in Appendix A.6). Common approaches to mitigate the reversal curse in autoregressive models include: (i) augmenting the training set with paraphrases (lu2024rethinking), which requires substantial computation to construct; (ii) augmenting the training set with reordered sequences (guo-etal-2024-mitigating; golovneva2024reverse), which often violate natural-language grammar and degrade overall language modeling performance; and (iii) replacing causal attention with bidirectional attention (lv-etal-2024-analysis), which struggles to retrieve long information (e.g., a person’s description). In contrast, our masked fine-tuning paradigm for arLLMs addresses the reversal curse without constructing paraphrase augmentations or altering the autoregressive objective. With only adjustments to the fine-tuning query and keeping the original documents as optimization targets, our method attains high accuracy across all datasets we evaluated.
@@ 2.2 Masked diffusion language models
Recently, dLLMs have emerged as a strong competitor to arLLMs (sahoo2024simple; llada; ye2025dream). Comparing to autoregressive models, dLLMs use encoder-only transformers to generate text by iteratively unmasking tokens via a reversed discrete diffusion process. The training objective is to minimize the mask reconstruction loss llada:
@@
$\bm{x_{0}}$ is an original sequence of length $L$ sampled from the training data. The masking process is governed by the mask ratio $t$, which is sampled uniformly, resulting in the corrupted sequence $\bm{x_{t}}$. The set $\mathbf{M}$ denotes the indices of the tokens that were masked by the forward process at ratio $t$. The $\ell$-th token is considered for the loss only if it was masked. Such a loss objective has been shown to be the negative evidence lower bound (ELBO) on the data likelihood (shi2024simplified).
Recent studies report dLLMs are more data-efficient than arLLMs. When the training data is scarce, dLLMs keep improving with repeated use of the data and surpass arLLMs on validation loss, while arLLMs saturate the validation loss or increase it due to overfitting (prabhudesai2025diffusion; ni2025difflm). prabhudesai2025diffusion further shows that the lower validation loss in dLLMs can generalize to downstream tasks like ARC-Easy, and attributes its data efficiency to random masks as implicit data augmentation. However, whether these advantages persist in new knowledge acquisition during post-training—where the model needs to learn generalizable knowledge through small fine-tuning sets—remains unclear.
Note that the denoising objective and bi-directional attention of dLLMs is closely related to the design choices in earlier language models such as BERT and T5/Flan-T5 devlin2019bert; raffel2020exploring. However, they are less popular choices for studying the post-training phase properties. BERT is not a generative chat models thus cannot achieve QA tasks; the complexity of the encoder-decoder design of T5 hinders its advancement to match other state-of-art LLMs. Thus, in this study, we choose to use state-of-the-art dLLMs and arLLMs to probe the post-training properties involving question-answering.
@@ 3 Datasets and experimental setups
We focus on assessing LLMs’ ability to learn new knowledge through fine-tuning. More specifically, LLMs are fine-tuned on a set of documents that contain knowledge unknown to the base LLM, and evaluated by open-ended QA tasks. The correctness of an answer is evaluated by the ROUGE-1 score (lin2024delving; jiang2025anyedit) between the generated answer and the ground truth answer, which we report as “accuracy.” It measures the proportion of the words in the ground truth answer that appear in the generated answer. To better demonstrate the generation quality, we also show examples of model responses in all the experiments in A.7.
@@
In the first part of the study, we use three representative datasets. Two are existing synthetic datasets from previous studies on the reversal curse, and each accompanied with a paraphrase set as data augmentation; one is constructed using recent Wikipedia articles. Previous studies on the reversal curse rarely use realistic datasets. In this study, we include the Wiki dataset to probe the reversal curse in a setting that reflects knowledge injection in the real world. See examples of each dataset in Appendix A.3.
@@
The NameDescription (berglund2023reversal) contains 60 statements of different fictitious individuals, 30 each of the form “[name] is [description]” (N2D) and “[description] is [name]” (D2N). lin2024delving extended the dataset with an open-ended QA testing set. For each type of statement, the QA set contains two types of questions: “What is the name related to a given description” and “What is the description of a given name”. Depending on whether the question is aligned with the original statement, each question is classified as “forward” or “backward” question (e.g. N2D statement with “What is the description of a given name” type of question is a forward question). The dataset also contains a paraphrase set, in which each statement is rewritten into 30 different versions, but the order of [name] and [description] in the paraphrases is always preserved as in the original statement (either N2D or D2N).
@@
The Biography dataset is proposed in zhu24a; zhu24b. Since the original dataset is not publicly available, we used a subset of 100 samples from a replication (zheng2025spurious). Each sample is a 6-sentence paragraph about a fictitious individual, detailing their birth city, birthday, college, and job information. Note that the name only appears in the first sentence and is replaced with a pronoun in the following sentences; thus, questions about the name are considered backward questions. Each sample also includes a paraphrase set of 5 paraphrases; the paraphrases do not change the order of the sentences but only alter the wording while preserving the information. The testing QA set has both forward (i.e., asking for an attribute given the name) and backward styles (i.e., asking for the name given 3 attributes from the person) questions.
@@
The Wiki dataset contains 94 Wikipedia articles, constructed according to the protocol described in pan2025memorization. We crawl the Wikipedia pages under the category “2025 by month”, and then further filter out the pages that were created before the year 2025. This procedure ensures that these real-world events are recent enough that both dLLM and arLLM models have minimal knowledge about them, which is justified by the models’ accuracy before fine-tuning (Table 1). For each wiki article, we use GPT-o3-mini to generate QA pairs in both forward and backward styles. By prompting GPT-o3-mini, we construct two different paraphrase sets: one that retains the information in place while only changing the wording (same-order paraphrases); the other also changes the order of information in the article (permute-order paraphrases). 10 paraphrases of each type are generated for every wiki article. All LLM generated QAs or paraphrases are cross-checked by human and invalid entries are filtered out and replaced. More details on constructing the datasets are provided in Appendix A.3.
@@
We chose the arLLMs Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct (dubey2024llama), Qwen2.5-7B-Instruct, Qwen3-4B-Instruct-2507 yang2025qwen3 and dLLM LLaDA-8B-Instruct (llada) to conduct the experiments. LLaDA-8B-Instruct is directly comparable to Llama-3.1-8B-Instruct as they have similar parameter size and comparable general capability benchmarks. Fine-tuning and evaluation configurations are provided in the Appendix A.4.
NameDescription Biography Wiki
N2D-fwd N2D-bwd D2N-fwd D2N-bwd Fwd Bwd Fwd Bwd
Llada before fine-tuning 0.030 0.000 0.028 0.000 0.030 0.000 0.210 0.156
Llada w/o paraphrases 0.873 0.913 0.864 0.790 0.892 0.696 0.908 0.778
Llada w paraphrases 0.967 0.994 0.994 0.973 0.991 0.857 0.900 0.785
Llama 8B before fine-tuning 0.072 0.000 0.054 0.000 0.001 0.000 0.164 0.127
Llama 8B reverse training 0.637 0.125 0.113 0.088 0.180 0.011 0.609 0.425
Llama 8B w/o paraphrases 0.374 0.000 0.017 0.027 0.121 0.002 0.377 0.282
Llama 8B w paraphrases 0.910 0.004 0.925 0.071 0.962 0.001 0.685 0.396
Masked Llama 8B w/o paraphrases 0.658 0.949 0.992 0.923 0.971 0.598 0.980 0.930
Masked Llama 8B w paraphrases 0.969 0.996 0.928 0.832 0.965 0.816 0.905 0.794
Llama 3B before fine-tuning 0.078 0.067 0.000 0.000 0.001 0.001 0.106 0.159
Llama 3B reverse training 0.391 0.079 0.042 0.025 0.216 0.003 0.384 0.285
Llama 3B w/o paraphrases 0.230 0.078 0.017 0.000 0.032 0.000 0.292 0.229
Llama 3B w paraphrases 0.951 0.040 0.967 0.025 0.988 0.001 0.622 0.334
Masked Llama 3B w/o paraphrases 0.887 0.932 0.992 0.933 0.967 0.738 0.970 0.908
Masked Llama 3B w paraphrases 0.982 0.928 1.000 0.992 0.964 0.809 0.855 0.810
Qwen 7B before fine-tuning 0.039 0.030 0.000 0.000 0.018 0.000 0.205 0.231
Qwen 7B reverse training 0.902 0.092 0.450 0.350 0.540 0.026 0.678 0.481
Qwen 7B w/o paraphrases 0.966 0.043 0.367 0.000 0.357 0.003 0.676 0.371
Qwen 7B w paraphrases 0.987 0.063 0.954 0.000 0.956 0.003 0.712 0.414
Masked Qwen 7B w/o paraphrases 0.870 0.897 0.967 0.929 0.984 0.754 0.934 0.896
Masked Qwen 7B w paraphrases 0.957 0.810 0.933 1.000 0.960 0.828 0.867 0.821
Qwen 4B before fine-tuning 0.026 0.032 0.000 0.000 0.058 0.001 0.237 0.243
Qwen 4B reverse training 0.259 0.083 0.008 0.008 0.140 0.008 0.470 0.437
Qwen 4B w/o paraphrases 0.137 0.081 0.000 0.000 0.103 0.001 0.509 0.353
Qwen 4B w paraphrases 0.962 0.038 0.975 0.000 0.994 0.003 0.675 0.418
Masked Qwen 4B w/o paraphrases 0.928 0.902 0.950 0.967 0.944 0.623 0.907 0.870
Masked Qwen 4B w paraphrases 0.975 0.816 1.000 0.975 0.967 0.806 0.847 0.822
Table 1: Fine-tuning performance of the dLLM and arLLMs across all three datasets. Model names are shortened for presentation (see section 3 for full model names). “Masked” denotes the masked fine-tuning paradigm. “Reverse training” denotes the entity-level reverse-training paradigm proposed in golovneva2024reverse, which we use as a baseline control. For the Wiki dataset, we use the same-order paraphrase set. Pink indicates clear failure (accuracy below 50%); turquoise indicates clear success (accuracy above 90%).
4 arLLM knowledge injection relies on paraphrases
We first show that knowledge injection by fine-tuning in arLLMs heavily relies on paraphrases. This is known in previous studies (berglund2023reversal; zhu24b; lin2024delving; guo-etal-2024-mitigating; golovneva2024reverse). We demonstrate this observation on three datasets to set baselines for comparison with dLLM and our novel paradigm in the following sections.
We fine-tune arLLMs on dataset samples using the pre-training format (i.e., without a chat template). Without paraphrases, backward accuracy on the NameDescription and Biography datasets is close to 0, while the forward accuracy of NameDescription N2D and Biography does not completely fail but is still poor (Table 1). Adding same-order paraphrases drastically raises forward accuracy close to 1, while backward accuracy remains close to 0. Paraphrases do not help backward accuracy in NameDescription and Biography datasets because the construction of these datasets does not change the semantic order of the sentences. The trend is similar in the Wiki dataset (Table 1 and 2). While the same-order paraphrases significantly increase forward accuracy, they only mildly increase backward accuracy. Using permute-order paraphrases increases both forward and backward accuracy, and the gap between them is smaller. Note that, due to the naturalness of this dataset, pre–fine-tuning accuracies are not as close to zero as in the other datasets (Table 1 and 2); nonetheless, they are sufficiently low to demonstrate the effectiveness of fine-tuning.
These results suggest that, in arLLM fine-tuning, paraphrases significantly improve QA accuracy, but help backward questions only when the paraphrases change the information order in the original text to be more aligned with the backward style. Note that the accuracy difference between fine-tuning with paraphrases and without paraphrases is not due to different training steps; in both cases, we train the models with sufficiently large epoch numbers; the reported accuracy is from the best checkpoints during the training (Figure 2, Appendix 6).
5 dLLM knowledge injection
@@
Llama 8B before fine-tuning 0.164 0.127
Llama 8B w/o paraphrases 0.377 0.282
Llama 8B w same-order paraphrases 0.685 0.396
Llama 8B w permute-order paraphrases 0.721 0.628
Llama 3B before fine-tuning 0.106 0.159
Llama 3B w/o paraphrases 0.292 0.229
Llama 3B w same-order paraphrases 0.622 0.334
Llama 3B w permute-order paraphrases 0.667 0.550
Qwen 7B before fine-tuning 0.205 0.231
Qwen 7B w/o paraphrases 0.676 0.371
Qwen 7B w same-order paraphrases 0.712 0.414
Qwen 7B w permute-order paraphrases 0.761 0.621
Qwen 4B before fine-tuning 0.237 0.243
Qwen 4B w/o paraphrases 0.509 0.353
Qwen 4B w same-order paraphrases 0.675 0.418
Qwen 4B w permute-order paraphrases 0.693 0.639
@@
Table 2: Fine-tuning performance of arLLM on the Wiki dataset. Pink indicates clear failure (accuracy below 50%).